### v1

@@ 5 Implementation and Benchmarks
We implemented the Shufflecake scheme in the C language as an open-source device-mapper-based driver for the Linux kernel. We published our code under the GPLv2+ license. The current release is v0.4.1 [45]. This section describes the programming environment and the structure of our implementation, and presents concrete performance measurements taking other popular disk encryption solutions as a baseline for comparison.
@@ 5.1 Structure of the Implementation
Our implementation consists of two components: a dm-sflc kernel module (which does most of the job), and a shufflecake companion userland application (used to correctly manage the volumes). The kernel module is the component that actually implements the scheme, translating logical requests into physical requests, and persisting slice maps into the respective headers.
@@
The other blocks of the volume header, which contain the slice map encrypted with the $\mathsf{VEK}$, are managed by the kernel module (except at init time, when an empty position map is written by the userland tool).
@@
The shufflecake init command takes as input a device path, and then interactively asks the user a number $\ell\leq\mathsf{max}$ and $\ell$ passwords as input, correctly formats the first $\ell$ volume headers, and fills the remaining $\mathsf{max}-\ell$ slots with random bytes; this way, the pre-existing volumes are wiped by erasing their headers (crypto-shredding). Unless a --skip-randfill option is provided (e.g., for testing or debugging purpose), the whole disk is filled with random bytes before formatting the header section. This command only formats the disk: it does not create the Linux virtual devices associated to the volumes.
The shufflecake open command takes a device path as input and asks one single password to the user, then looks up the volume headers, and opens all the volumes starting from the one whose password is provided, backwards up to the first one (walking up the chain using the $\mathsf{VMK}_{i-1}$ field in the VMB). This is the command that actually creates the Linux virtual devices representing the volumes, under /dev/mapper: the names are generated algorithmically. Notice that these virtual devices are not automatically mounted, it is up to the user to mount them and format them with a filesystem of choice when required.
Finally, the shufflecake close command takes a device path as input, and closes all the volumes open on that device.
@@ 5.2 Space Utilisation
For our implementation, we fixed the block size to 4096 bytes, so as to better amortise the per-block space overhead determined by the IVs. We chose $S_{L}=256$, and $\mathsf{max}=15$. Since we use AES-CTR-256 as the underlying encryption scheme, we need 16-byte IVs. This led to a choice of $\Delta_{S}=1$: a single 4-KiB IV block (containing 256 IVs) encrypts a 1-MiB slice. To provide a numerical summary of the space utilisation of Shufflecake, we observe that in the case of a 1 TiB disk, the resulting theoretical maximum utilisable space is 1019.91 GiB, equal to more than 99.6% of the physical storage space.
@@
With the above parameters, the total size of a volume header is around $\frac{N}{S_{P}}\log\frac{N}{S_{P}}$, roughly equal to 4 MiB per volume header, for a 1-TiB disk: about 60 MiB for the total device header size.
@@
Internal fragmentation is a frequent problem in space allocation, and it is particularly well known and studied in file systems theory. For performance reasons, the block layer only works with the block granularity; the file system, therefore, has to allocate a whole block even if it needs less space to, e.g., host a file. Internal fragmentation is the problem arising from this “over-allocation”. On top of this, Shufflecake adds another layer of internal fragmentation through its slice mechanism: when a volume requests a block, we reserve a whole slice of $S_{L}$ blocks just for that volume. Moreover, we have no means of communicating this over-allocation to the file system layer, which therefore has no way of adapting its behaviour. Thus, we have to hope that a file system does not jump back and forth too wildly, and that it generally tries to fill a group of slices before requesting a new one.
Luckily, some file systems do exhibit this behaviour. For example, the commonly used ext4 file system defines the concept of a block group, i.e., a group consisting of 32768 consecutive blocks (which amounts to 128 MiB for 4096-byte blocks). The block allocator of ext4 tries its hardest to keep related files within the same block group; specifically, whenever possible, it stores all inodes of a directory in the same block group as the directory; also, it stores all blocks of a file in the same block group as its inode [27]. This feature plays nicely with the value of $S_{L}=256$ we chose: a block group encompasses a whole number of slices, which will therefore not be too fragmented in the long run.
@@ 5.3 Benchmarks
We tested our implementation looking at I/O performance and space efficiency. The test environment was a fresh installation of Ubuntu 23.04 running kernel version 6.2.0 on a laptop equipped with an AMD Ryzen 7 PRO 6850U CPU with Spectre mitigations enabled, 32 GiB 4-channel 6400 MHz DDR5 RAM, and a low-power 1 TiB NVMe Micron MTFDKCD1T0TFK SSD. We tested the amount of slice fragmentation of Shufflecake (v0.4.1), its I/O performance, as well as the I/O performance of other two relevant disk-encryption tools for comparison: dm-crypt/LUKS (v6.2.0-26) and VeraCrypt (v1.25.9). All the tests were performed sequentially, on a physical primary SSD partition of size 8 GB, using the ext4 filesystem (which is the one most relevant for Shufflecake’s envisioned final use case). In the case of Shufflecake, we initiated the partition with two volumes (one decoy and one hidden), and performed all tests on the hidden one. Analogously, in the case of VeraCrypt, we formatted the partition as a standard no-FS VeraCrypt volume, and created a 6.5 GB ext4 volume therein. In order to aid reproducibility, we also included in our implementation a suite of benchmark scripts performing the tests described here.
@@ 5.3.1 Fragmentation
In order to evaluate the fragmentation caused by Shufflecake’s allocation of slices, we filled the ext4 filesystem with incrementally large amount of random files and directories up to saturating the space, and at every step we measured the space given by the increasing number of slices allocated by dm-sflc for the hidden volume VS the total amount of data written therein. We define space efficiency as the ratio between real data written on disk and slice-allocated space (00 = bad, $1$ = good).
[ xlabel=Volume occupancy ratio, ylabel=Space efficiency, xmin=0, xmax=0.95, ymin=0, ymax=1.0, xtick=0,.1,.2,.3,.4,.5,.6,.7,.8,.9, ytick=0,.2,.4,.6,.8,1.0, legend pos=north west, ymajorgrids=true, grid style=dashed, ]
[ color=blue, mark=dot, thick, ] coordinates (0,0) (0.05,0.8197) (0.1,0.9009) (0.15,0.9317) (0.2,0.9479) (0.25,0.9579) (0.3,0.9677) (0.35,0.9695) (0.4,0.9732) (0.45,0.9719) (0.5,0.9728) (0.55,0.9717) (0.6,0.9724) (0.65,0.9716) (0.7,0.9709) (0.75,0.9715) (0.8,0.9721) (0.85,0.9714) (0.9,0.9730) (0.95,0.9744) ;
@@
The comparison with some popular ORAM-based PD solutions more convincingly shows the real-world efficiency advantages offered by Shufflecake. Of course it has to be stressed that these ORAM-based solutions aim at achieving PD in a more strict scenario than the single-snapshot security offered by the current version of Shufflecake. We have just seen how Shufflecake achieves a slowdown of roughly 30% I/O throughput over dm-crypt and uses almost all space available. On the other hand, HiVE [6] has a heavy 200x I/O overhead and wastes 50% of the disk space. DetWoOram [39] has an overhead of 2.5x for reads and 10x-14x for writes, and wastes 75% of space.
@@ 6.1 Crash Consistency
An alternative solution, which wastes more disk space but we believe to be overall better, would be to store the IV alongside the data block itself, so that the two updates can be merged into a single physical request. We believe that, in addition to mitigate the issue of crash inconsistency, this approach would probably lead to better I/O performance, as it would not need separate operations for writing IV blocks during data writes. The minimum addressable unit of disk storage space, at least on Linux systems, is usually the 512-byte sector. Therefore, the least wasteful option is to map a logical 4096-byte block (8 sectors, as was already the case for our implementation) onto 9 consecutive physical sectors: the first one contains the IV, the other ones constitute the data block. This would lead to a waste of disk space (fraction of disk not used for upper-layer data) equal to $\frac{1}{9}=11.1\%$. Since an IV only occupies 16 bytes, much of the first sector would be left unused; we could use the rest of the free space to also contain additional useful data, for example a MAC to detect IV alterations, and a “reverse map”, indicating which logical block $B$ of which volume $V_{i}$ that physical block corresponds to (this information would of course be encrypted). Additionally, we could build an even more fault-resilient system by again having a circular log of two IVs in the first sector (each accompanied by the corresponding MAC), plus 16 MACs, one for each IV and for each of the 8 data sectors. This would allow us to disambiguate with the sector granularity, in case the underlying disk can ensure atomicity of the sector writes, but not of the physical requests writing on several adjacent sectors.
@@ 6.2 Multi-Snapshot Security
Physical slices have known boundaries, so the adversary can compare the disk snapshots she has on a slice basis. This amounts to inspecting the subsequent changes, or diffs, each physical slice goes through across snapshots. Recall that a physical slice is essentially an array of $S_{P}$ blocks; when one of these blocks changes, it can either be because of the re-encryption of a data block (with a different IV, and possibly a different content), or because of the re-randomisation of an empty block: the IND-CPA security of the encryption scheme guarantees the indistinguishability of these two situations. This means that no hint on the nature of a block (i.e., whether it is a data or an empty block) is leaked by its encrypted content or by the history of its encrypted contents. Therefore, when comparing two snapshots of a physical slice, the only information that an adversary gets is which of the $S_{P}$ blocks have changed: how they have changed is completely inconsequential and uninformative. In other words, the diff between two snapshots of a physical slice boils down to a bitmask of $S_{P}$ bits, representing which blocks have changed and which remained the same.
@@
The task of the adversary then becomes to distinguish between “data slices” (i.e., belonging to some volume) and “free slices” (not mapped to any volume) based on a sequence of such diffs for each slice. The point is that, after unlocking the first $\ell-1$ volumes, the rest of the space may or may not contain another volume (there may or may not be some more data slices among the free slices): if we want the adversary to be incapable of distinguishing between the two cases, we need the diffs of data slices and free slices to look the same. Once we frame the problem in this way, we can rephrase the weakness of Shufflecake as follows: the diff bitmask of a free slice is always all-zeros, and as such is clearly distinguishable from that of a data slice, which might have some bits set to $1$.
@@
Our task then becomes to “obfuscate” the changes that occurred in the data slices (especially those belonging to $V_{\ell}$) by artificially creating a non-zero diff bitmask in the free slices, through a re-randomisation of some selected blocks. This way, a user will hopefully be able to claim that all the changes happened to the “empty” blocks are due to this obfuscation procedure, and not to the existence of $V_{\ell}$. Nothing prevents us, of course, from also touching the data slices during the obfuscation procedure, if that helps making the diff bitmasks look more alike. It is to be noted, however, that if a block was modified by the upper layer during the normal operational phase, the corresponding bit will be set to 1 in the diff bitmask and there is no way for the obfuscation procedure to “undo” that change: when we touch a data slice, we cannot turn the $1$s of the diff bitmask into 00s. Instead, we can turn some 00s into 1s by simply re-encrypting the same content of a block with a different IV.
@@
A very simple first idea for an obfuscation procedure is to take all physical blocks belonging to free slices, and re-randomise them all, independently at random, each with probability $p$. This operation could be either performed upon volume close or, for better resilience, spread across the normal operations. It is the easiest way to achieve a non-zero diff bitmask for free slices, but it is definitely too crude to work: nothing guarantees that the diff bitmasks of data slices will “look random” like the ones artificially generated for free slices. Also, it might very well be the case that many data slices do not change across two snapshots, in which case it becomes very easy to tell them apart from free slices, which often have a non-zero diff bitmask.
Such consideration suggests a refinement: besides re-randomising some blocks in the free slices, we could re-encrypt (with a different IV but same plaintext) some blocks in the data slices of existing volumes, again independently at random, each with probability $q$. This way, we also randomise the diff bitmasks of the data slices, making them more similar to those generated for free slices.
@@
The procedure we just illustrated could well succeed, for a suitable choice of $p$ and $q$, in rendering the diff bitmasks of data slices and free slices roughly indistinguishable, but only if the adversary gets just one diff for each slice, i.e., if she only gets two snapshots. This is because we are essentially playing a hopeless game: roughly speaking, we are aiming at making signal+noise (the diffs of a data slice) indistinguishable from noise alone (the diff of a free slice). As discussed, we cannot turn the $1$s of the diff bitmasks of data slices into 00s: the “signal” given by which blocks were modified by the upper layer stays there, we can only hope to bury it in enough noise by turning some 00s into $1$s through re-encryption. However, with enough snapshots, the signal will eventually emerge. Imagine, for instance, that there is one particular block in a data slice that is very often modified (maybe it contains some sort of file system index): the corresponding bit in the diff bitmask will often be set to $1$, which would be hard to justify through the obfuscation procedure, which only hits one given block with probability $p$ each time.
@@
The previous discussion teaches us a valuable lesson: in our setting, we cannot hope to disguise the accesses performed by the upper layer as random noise, as is commonly the case for ORAMs. The only option we have left is to take the opposite approach: let us make the diffs of free slices look like they were also generated by some file system workload. This way, we are still making the diffs for the two kinds of slices similar, but we are not trying to erase the signal from existing data slices, which we cannot. Instead, we “copy this signal” onto the free slices, so that the adversary will always see this signal, even if the last $\ell$-th volume has been surrendered. A tedious, convoluted, and yet very imprecise way of doing it would be for us to sit down, study the access patterns resulting from typical file system workloads, model them as a probability distribution, and hardcode that into the obfuscation procedure. Probably, a better idea would be to use a ML approach and let a daemon run in the background to adaptively learn and simulate such distribution.
Instead, a simpler method that is likely to capture the patterns we want to imitate is to have the scheme itself “learn” them online, by simply subsampling the stream of incoming logical requests. More concretely: for each incoming logical SflcWrite request, we “imitate” it with probability $p$. Imitating a request means “learning” that the affected logical block $b$ is likely to be written by file system workloads, and copy this signal onto a free slice: we retain the offset $(b\mod S_{L})$ of the block within the slice, we choose a “target” free slice, and we re-randomise the block with the same offset within the target slice. This approach guarantees that, if a particular block is often updated by the upper layer, then we are very likely to catch this signal and correctly carry it over to a free slice. Note, however, that we need to choose the target free slice deterministically from some context information. Only in this way can we replicate the signal consistently across snapshots, always onto the same free slice; also, this allows us to correctly capture and copy a signal in case it consists of not just one, but several blocks in a slice being frequently updated.
@@
If we assume that the obfuscation procedure just described really succeeds in making the empty space look like it is occupied, we are still left with one problem. The special blocks that are often updated by the upper layer leave a very clear trace in the snapshot history, since their bits in the diff bitmasks are almost always set to 1. If we have $\ell$ volumes, the obfuscation procedure will generate $\ell$ such clear traces in the free space. Therefore, a simple attack would consist in counting these traces, and checking whether there are as many of them as there are disclosed volumes. To thwart this attack, we can rework the obfuscation procedure in such a way that the device’s data section always looks like it’s hosting $\mathtt{max}$ volumes. We can ideally assign the free slices to $\mathtt{max}-\ell$ pairwise-disjoint sets, each one representing a “fake” volume, and have the obfuscation procedure be aware of this partitioning when choosing the target free slice, so as to really simulate $\mathtt{max}-\ell$ volumes with the imitated logical SflcWrite requests.
@@
The obfuscation procedure described above may already offer good protection, although it might be non-trivial to translate the rough idea of “being aware of the partitioning into $\mathtt{max}-\ell$ fake volumes” into a concrete algorithm for choosing a target free slice when we imitate a SflcWrite request. A valid, if “exotic”, alternative, would be to actually create $\mathtt{max}-\ell$ additional “ghost” Shufflecake volumes on the device, that behave in the exact same way as regular volumes, except that they do give up their slices when needed. On these volumes, a separate component (a daemon) could mount a file system and perform some typical sequences of accesses. The advantage of this solution is that, by definition, it will always look like there are $\mathtt{max}$ volumes on the device. However, when slices are reclaimed from ghost volumes, their file systems might suddenly get corrupted and start complaining quite loudly, thus impairing the practical usability of the system. A daemon operating on these ghost filesystems might therefore need to be aware of the new slice allocation requests from “real” volumes, and move or remap the ghost slices accordingly at runtime.

### v2

@@ 5 Implementation and Benchmarks
We implemented the Shufflecake scheme in the C language as an open-source device-mapper-based driver for the Linux kernel. We published our code under the GPLv2+ license. The current release is v0.4.3 [45]. This section describes the programming environment and the structure of our implementation, and presents concrete performance measurements taking other popular disk encryption solutions as a baseline for comparison.
@@ 5.1 Structure of the Implementation
Our implementation consists of two components: a $\mathtt{dm\mbox{-}sflc}$ kernel module (which does most of the job), and a $\mathtt{shufflecake}$ companion userland application (used to correctly manage the volumes). The kernel module is the component that actually implements the scheme, translating logical requests into physical requests, and persisting slice maps into the respective headers.
@@
The other blocks of the volume header, which contain the slice map encrypted with the $\mathsf{VEK}$, are managed by the kernel module (except at $\mathtt{init}$ time, when an empty slice map is written by the userland tool).
@@
The $\mathtt{shufflecake}$ $\mathtt{init}$ command takes as input a device path, and then interactively asks the user a number $N\leq N_{\mathsf{MAX}}$ and $N$ passwords as input, correctly formats the first $N$ volume headers, and fills the remaining $N_{\mathsf{MAX}}-N$ slots with random bytes; this way, the pre-existing volumes are wiped by erasing their headers (crypto-shredding). Unless a --skip-randfill option is provided (e.g., for testing or debugging purpose), the whole disk is filled with random bytes before formatting the header section. This command only formats the disk: it does not create the Linux virtual devices associated to the volumes.
The $\mathtt{shufflecake}$ $\mathtt{open}$ command takes a device path as input and asks one single password to the user, then looks up the volume headers, and opens all the volumes starting from the one whose password is provided, backwards up to the first one (walking up the chain using the $\mathsf{VMK}_{i-1}$ field in the VMB). This is the command that actually creates the Linux virtual devices representing the volumes, under /dev/mapper: the names are generated algorithmically. Notice that these virtual devices are not automatically mounted, it is up to the user to mount them and format them with a filesystem of choice when required.
Finally, the $\mathtt{shufflecake}$ $\mathtt{close}$ command takes a device path as input, and closes all the volumes open on that device.
@@ 5.2 Space Utilisation
For our implementation, we fixed the block size $S_{\mathsf{B}}$ to 4096 bytes (32768 bits), so as to better amortise the per-block space overhead determined by the IVs. We chose $S_{\mathsf{L}}=256$, and $N_{\mathsf{MAX}}=15$. Since we use AES-CTR-256 as the underlying encryption scheme, we need 16-byte IVs. This led to a choice of $\Delta_{\mathsf{S}}=1$, and hence $S_{\mathsf{P}}=257$: a single 4-KiB IV block (containing 256 IVs) encrypts a 1-MiB slice. To provide a numerical summary of the space utilisation, we observe that in the case of a 1 TiB disk, the resulting theoretical maximum utilisable space is 1019.91 GiB, equal to more than 99.6% of the physical storage space.
@@
With the above parameters, the total size of a volume header is around $S_{\mathsf{H}}=\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}}\log\frac{S_{\mathsf{D}}}{S_{\mathsf{P}}}$, roughly equal to 4 MiB per volume header, for a 1-TiB disk: about 60 MiB for the total device header size.
@@
Internal fragmentation is a frequent problem in space allocation, and it is particularly well known and studied in file systems theory. For performance reasons, the block layer only works with the block granularity; the file system, therefore, has to allocate a whole block even if it needs less space to, e.g., host a file. Internal fragmentation is the problem arising from this “over-allocation”. On top of this, Shufflecake adds another layer of internal fragmentation through its slice mechanism: when a volume requests a block, we reserve a whole slice of $S_{\mathsf{L}}$ blocks just for that volume. Moreover, we have no means of communicating this over-allocation to the file system layer, which therefore has no way of adapting its behaviour. Thus, we have to hope that a file system does not jump back and forth too wildly, and that it generally tries to fill a group of slices before requesting a new one.
Luckily, some file systems do exhibit this behaviour. For example, the commonly used ext4 file system defines the concept of a block group, i.e., a group consisting of 32768 consecutive blocks (which amounts to 128 MiB for 4096-byte blocks). The block allocator of ext4 tries its hardest to keep related files within the same block group; specifically, whenever possible, it stores all inodes of a directory in the same block group as the directory; also, it stores all blocks of a file in the same block group as its inode [27]. This feature plays nicely with the value of $S_{\mathsf{L}}=256$ we chose: a block group encompasses a whole number of slices, which will therefore not be too fragmented in the long run.
@@ 5.3 Benchmarks
We tested our implementation looking at I/O performance and space efficiency. The test environment was a fresh installation of Ubuntu 23.04 running kernel v6.2.0 on a laptop equipped with an AMD Ryzen 7 PRO 6850U CPU with Spectre mitigations enabled, 32 GiB 4-channel 6400 MHz DDR5 RAM, and a low-power 1 TiB NVMe Micron MTFDKCD1T0TFK SSD. We tested the amount of slice fragmentation of Shufflecake (v0.4.1), its I/O performance, as well as the I/O performance of other two relevant disk-encryption tools for comparison: dm-crypt/LUKS (v6.2.0-26) and VeraCrypt (v1.25.9). All the tests were performed sequentially, on a physical primary SSD partition of size 8 GB, using the ext4 filesystem (which is the one most relevant for Shufflecake’s envisioned final use case). In the case of Shufflecake, we $\mathtt{init}$iated the partition with two volumes (one decoy and one hidden), and performed all tests on the hidden one. Analogously, in the case of VeraCrypt, we formatted the partition as a standard no-FS VeraCrypt volume, and created a 6.5 GB ext4 volume therein. In order to aid reproducibility, we also included in our implementation a suite of benchmark scripts performing the tests described here.
@@ 5.3.1 Fragmentation
In order to evaluate the fragmentation caused by Shufflecake’s allocation of slices, we filled the ext4 filesystem with incrementally large amount of random files and directories up to saturating the space, and at every step we measured the space given by the increasing number of slices allocated by $\mathtt{dm\mbox{-}sflc}$ for the hidden volume VS the total amount of data written therein. We define space efficiency as the ratio between real data written on disk and slice-allocated space (00 = bad, $1$ = good).
000.10.10.10.10.20.20.20.20.30.30.30.30.40.40.40.40.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.80.90.90.90.9000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111Volume occupancy ratioSpace efficiency
@@
The comparison with some popular ORAM-based PD solutions more convincingly shows the real-world efficiency advantages offered by Shufflecake. Of course it has to be stressed that these ORAM-based solutions aim at achieving PD in a more strict scenario than the single-snapshot security offered by the current version of Shufflecake. We have just seen how Shufflecake achieves a slowdown of roughly 30% I/O throughput over dm-crypt and uses almost all space available. On the other hand, HiVE [6] has a heavy 200x I/O overhead and wastes 50% of the disk space. DetWoOram [39] has an overhead of 2.5x for $\mathtt{read}$s and 10x-14x for $\mathtt{write}$s, and wastes 75% of space.
@@ 6.1 Crash Consistency
An alternative solution, which wastes more disk space but we believe to be overall better, would be to store the IV alongside the data block itself, so that the two updates can be merged into a single physical request. We believe that, in addition to mitigate the issue of crash inconsistency, this approach would probably lead to better I/O performance, as it would not need separate operations for writing IV blocks during data writes. The minimum addressable unit of disk storage space, at least on Linux systems, is usually the 512-byte sector. Therefore, the least wasteful option is to map a logical 4096-byte block (8 sectors, as was already the case for our implementation) onto 9 consecutive physical sectors: the first one contains the IV, the other ones constitute the data block. This would lead to a waste of disk space (fraction of disk not used for upper-layer data) equal to $\frac{1}{9}=11.1\%$. Since an IV only occupies 16 bytes, much of the first sector would be left unused; we could use the rest of the free space to also contain additional useful data, for example a MAC to detect IV alterations, and a “reverse map”, indicating which logical block $\beta_{\mathsf{L}}$ of which volume $V_{i}$ that physical block corresponds to (this information would of course be encrypted). Additionally, we could build an even more fault-resilient system by again having a circular log of two IVs in the first sector (each accompanied by the corresponding MAC), plus 16 MACs, one for each IV and for each of the 8 data sectors. This would allow us to disambiguate with the sector granularity, in case the underlying disk can ensure atomicity of the sector writes, but not of the physical requests writing on several adjacent sectors.
@@ 6.2 Multi-Snapshot Security
Physical slices have known boundaries, so the adversary can compare the disk snapshots she has on a slice basis. This amounts to inspecting the subsequent changes, or diffs, each physical slice goes through across snapshots. Recall that a physical slice is essentially an array of $S_{\mathsf{P}}$ blocks; when one of these blocks changes, it can either be because of the re-encryption of a data block (with a different IV, and possibly a different content), or because of the re-randomisation of an empty block: the IND-CPA security of the encryption scheme guarantees the indistinguishability of these two situations. This means that no hint on the nature of a block (i.e., whether it is a data or an empty block) is leaked by its encrypted content or by the history of its encrypted contents. Therefore, when comparing two snapshots of a physical slice, the only information that an adversary gets is which of the $S_{\mathsf{P}}$ blocks have changed: how they have changed is completely inconsequential and uninformative. In other words, the diff between two snapshots of a physical slice boils down to a bitmask of $S_{\mathsf{P}}$ bits, representing which blocks have changed and which remained the same.
@@
The task of the adversary then becomes to distinguish between “data slices” (i.e., belonging to some volume) and “free slices” (not mapped to any volume) based on a sequence of such diffs for each slice. The point is that, after unlocking the first $N-$ volumes, the rest of the space may or may not contain another volume (there may or may not be some more data slices among the free slices): if we want the adversary to be incapable of distinguishing between the two cases, we need the diffs of data slices and free slices to look the same. Once we frame the problem in this way, we can rephrase the weakness of Shufflecake as follows: the diff bitmask of a free slice is always all-zeros, and as such is clearly distinguishable from that of a data slice, which might have some bits set to $1$.
@@
Our task then becomes to “obfuscate” the changes that occurred in the data slices (especially those belonging to $V_{N-1}$) by artificially creating a non-zero diff bitmask in the free slices, through a re-randomisation of some selected blocks. This way, a user will hopefully be able to claim that all the changes happened to the “empty” blocks are due to this obfuscation procedure, and not to the existence of $V_{N-1}$. Nothing prevents us, of course, from also touching the data slices during the obfuscation procedure, if that helps making the diff bitmasks look more alike. It is to be noted, however, that if a block was modified by the upper layer during the normal operational phase, the corresponding bit will be set to 1 in the diff bitmask and there is no way for the obfuscation procedure to “undo” that change: when we touch a data slice, we cannot turn the $1$s of the diff bitmask into 00s. Instead, we can turn some 00s into 1s by simply re-encrypting the same content of a block with a different IV.
@@
A very simple first idea for an obfuscation procedure is to take all physical blocks belonging to free slices, and re-randomise them all, independently at random, each with probability $p_{1}$. This operation could be either performed upon volume $\mathtt{close}$ or, for better resilience, spread across the normal operations. It is the easiest way to achieve a non-zero diff bitmask for free slices, but it is definitely too crude to work: nothing guarantees that the diff bitmasks of data slices will “look random” like the ones artificially generated for free slices. Also, it might very well be the case that many data slices do not change across two snapshots, in which case it becomes very easy to tell them apart from free slices, which often have a non-zero diff bitmask.
Such consideration suggests a refinement: besides re-randomising some blocks in the free slices, we could re-encrypt (with a different IV but same plaintext) some blocks in the data slices of existing volumes, again independently at random, each with probability $p_{2}$. This way, we also randomise the diff bitmasks of the data slices, making them more similar to those generated for free slices.
@@
The procedure we just illustrated could well succeed, for a suitable choice of $p_{1}$ and $p_{2}$, in rendering the diff bitmasks of data slices and free slices roughly indistinguishable, but only if the adversary gets just one diff for each slice, i.e., if she only gets two snapshots. This is because we are essentially playing a hopeless game: roughly speaking, we are aiming at making signal+noise (the diffs of a data slice) indistinguishable from noise alone (the diff of a free slice). As discussed, we cannot turn the $1$s of the diff bitmasks of data slices into 00s: the “signal” given by which blocks were modified by the upper layer stays there, we can only hope to bury it in enough noise by turning some 00s into $1$s through re-encryption. However, with enough snapshots, the signal will eventually emerge. Imagine, for instance, that there is one particular block in a data slice that is very often modified (maybe it contains some sort of file system index): the corresponding bit in the diff bitmask will often be set to $1$, which would be hard to justify through the obfuscation procedure, which only hits one given block with probability $p_{1}$ each time.
@@
The previous discussion teaches us a valuable lesson: in our setting, we cannot hope to disguise the accesses performed by the upper layer as random noise, as is commonly the case for ORAMs. The only option we have left is to take the opposite approach: let us make the diffs of free slices look like they were also generated by some file system workload. This way, we are still making the diffs for the two kinds of slices similar, but we are not trying to erase the signal from existing data slices, which we cannot. Instead, we “copy this signal” onto the free slices, so that the adversary will always see this signal, even if the last $N$-th volume has been surrendered. A tedious, convoluted, and yet very imprecise way of doing it would be for us to sit down, study the access patterns resulting from typical file system workloads, model them as a probability distribution, and hardcode that into the obfuscation procedure. Probably, a better idea would be to use a ML approach and let a daemon run in the background to adaptively learn and simulate such distribution.
Instead, a simpler method that is likely to capture the patterns we want to imitate is to have the scheme itself “learn” them online, by simply subsampling the stream of incoming logical requests. More concretely: for each incoming logical $\mathtt{SflcWrite}$ request, we “imitate” it with probability $p$. Imitating a request means “learning” that the affected logical block at position $\beta_{\mathsf{L}}$ is likely to be written by file system workloads, and copy this signal onto a free slice: we retain the offset $(\beta_{\mathsf{L}}\mod S_{\mathsf{L}})$ of the block within the slice, we choose a “target” free slice, and we re-randomise the block with the same offset within the target slice. This approach guarantees that, if a particular block is often updated by the upper layer, then we are very likely to catch this signal and correctly carry it over to a free slice. Note, however, that we need to choose the target free slice deterministically from some context information. Only in this way can we replicate the signal consistently across snapshots, always onto the same free slice; also, this allows us to correctly capture and copy a signal in case it consists of not just one, but several blocks in a slice being frequently updated.
@@
If we assume that the obfuscation procedure just described really succeeds in making the empty space look like it is occupied, we are still left with one problem. The special blocks that are often updated by the upper layer leave a very clear trace in the snapshot history, since their bits in the diff bitmasks are almost always set to 1. If we have $N$ volumes, the obfuscation procedure will generate $N$ such clear traces in the free space. Therefore, a simple attack would consist in counting these traces, and checking whether there are as many of them as there are disclosed volumes. To thwart this attack, we can rework the obfuscation procedure in such a way that the device’s data section always looks like it’s hosting $N_{\mathsf{MAX}}$ volumes. We can ideally assign the free slices to $N_{\mathsf{MAX}}-N$ pairwise-disjoint sets, each one representing a “fake” volume, and have the obfuscation procedure be aware of this partitioning when choosing the target free slice, so as to really simulate $N_{\mathsf{MAX}}-N$ volumes with the imitated logical $\mathtt{SflcWrite}$ requests.
@@
The above procedure may already offer good protection, although it might be non-trivial to translate the rough idea of “being aware of the partitioning into $N_{\mathsf{MAX}}-N$ fake volumes” into a concrete algorithm for choosing a target free slice when we imitate a $\mathtt{SflcWrite}$ request. A valid, if “exotic”, alternative, would be to actually create $N_{\mathsf{MAX}}-N$ additional “ghost” Shufflecake volumes on the device, that behave in the exact same way as regular volumes, except that they do give up their slices when needed. On these volumes, a separate component (a daemon) could mount a file system and perform some typical sequences of accesses. The advantage of this solution is that, by definition, it will always look like there are $N_{\mathsf{MAX}}$ volumes on the device. However, when slices are reclaimed from ghost volumes, their file systems might suddenly get corrupted and start complaining quite loudly, thus impairing the practical usability of the system. A daemon operating on these ghost filesystems might therefore need to be aware of the new slice allocation requests from “real” volumes, and move or remap the ghost slices accordingly at runtime.